A tutorial class contains 7 male students and 5 female students. Find the number n of ways that the class can elect:
a) a class representative.
b) two class representatives, one male and one female.
c) a class representative and a deputy class representative.

Answer by ikleyn(52778)

.

All these questions are quite elementary.

    (a)  7+5   =  12   ways.


    (b)  7*5   =  35   ways.


    (c)  12*11 = 132   ways.

These answers are SELF-EVIDENT and do not require explanations.